OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $24.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount?

OpenStudy (anonymous):

\[24+.06x=30+.05x\] solve for x maybe easier if you write \[2400+6x=3000+5x\]

OpenStudy (anonymous):

\[6x-5x=3000-2400\] \[x=600\]

sam (.sam.):

Simple algebra problem, see below for the work. 24 +0.06*x = 30 + .05x Subtract 24 from both sides and subtract 0.05x from both sides 0.01x = 6 Divide 6 by 0.01 x = 600

OpenStudy (anonymous):

Cell Plus monthly cost P: 24 + .06m Cell Best monthly cost B: 30 + .05m P = B 24 + .06m = 30 + .05m .01m = 30 -24 = 6 =>m = 6/ .01 = 600 min